OpenStudy (anonymous):

Find dy/dx. y=(sqrt{x}-1)/(sqrt{x}+1)

OpenStudy (anonymous):

Can you show me the steps or tell me what I am doing wrong?

OpenStudy (valpey):

\[v = x^{1/2} - 1:: dv = (1/2)x^{-1/2}\]

OpenStudy (valpey):

Similarly if \[u = x^{1/2}+1 :: du = (1/2)x^{-1/2}\]

OpenStudy (lalaly):

use the quotient rule \[y=\frac{\sqrt x-1}{\sqrt x+1}\] \[f(x)=\sqrt x-1\]\[g(x)=\sqrt x+1\] \[\frac{dy}{dx}= \frac{f'(x)g(x)-f(x)g'(x)}{(g(x))^2}\]

OpenStudy (anonymous):

Yes.. I believe that is what I have? I'm not sure how to simply/condense it.

OpenStudy (valpey):

Oops I have my u and v backward. But it looked like you missed the exponent = -1/2 in your derivatives.

OpenStudy (anonymous):

Ohh okay! Thanks :) But I am still not sure how to simplify it.  Will it be something like this?